OpenStudy (anonymous):

what are the solutions of 16x^2 + 9 = 0?

OpenStudy (anonymous):

subtract 9 on both sides do it i will help u:)

OpenStudy (anonymous):

so take 16x^2 - 9?

OpenStudy (anonymous):

no subtract both sides -9

OpenStudy (anonymous):

ok so 9-9?

OpenStudy (anonymous):

yep this in the left side i the right side 0-9=-9

OpenStudy (anonymous):

then would I divide 16x^2 by -9?

OpenStudy (anonymous):

no divide 16 on both sides

OpenStudy (anonymous):

16x/16x = 0 then 16/-9?

OpenStudy (anonymous):

x^2=-9/16

OpenStudy (anonymous):

r u still there coz i hv to go?

OpenStudy (anonymous):

thanks for your help (: